OpenStudy (anonymous):

(4x)^2(x) Can someone help me solve this?

OpenStudy (ash2326):

Is this your question brad? \[(4x)^2 \times x\]

OpenStudy (anonymous):

yes\

OpenStudy (ash2326):

Ok, \[ (4x)^2 \times x\] let's open the bracket \[4^2 \times x^2\times x\] we know that 4^2=16 so \[16x^2\times x\] and we know that \(x^2\times x=x^3\) so we get \[16x^3\]

OpenStudy (anonymous):

can you help me with one more problem?

OpenStudy (ash2326):

Yeah surely post it as a new question. Welcome to Open Study:)

OpenStudy (anonymous):

how do i post a new question im very new lol

OpenStudy (ash2326):

On the left hand side, you'll see written "Ask a question", just type your question there